OpenStudy (anonymous):

PLEASE HELP me with this? :) Assume a 115-pound person burns 36 calories after 12 minutes of walking and 54 calories after 18 minutes of walking. Part 1: Use calories as the y-coordinate and minutes as the x-coordinate. After writing two ordered pairs, find the slope. What does the slope represent in terms of the information provided? Part 2 : Write an equation, in slope intercept form, to represent this data. Part 3: How many calories will this person burn after 35 minutes of walking? Using complete sentences, explain how the equation, slope, or graph can help to predict calories burned.

OpenStudy (anonymous):

Pleasee helpp

OpenStudy (anonymous):

We use the two point formula of straight line: \[y-y1=((y2-y1)\div (x2-x1)) \times (x - x1)\] Substituting for y1=36, x1 = 12, y2 = 54, x2 = 18, we get the equation as: \[y=3x\] Now substitute x=35 and you will get y as the calories. \[y = 3 \times 35\]\[y = 105\]